OpenStudy (anonymous):

i cant solve this--->x^3y^3 - 4=0...find d2y/dx2 by implicit differentiation..please....

OpenStudy (anonymous):

You have to use the product rule for the first term.

OpenStudy (anonymous):

\[\frac{d}{dx}[x^3y^3] = x^3\frac{d}{dx}[y^3] + y^3\frac{d}{dx}[x^3]\]

OpenStudy (anonymous):

Then recall that \[\frac{d}{dx}f(y) = \frac{d}{dy}f(y)\cdot \frac{dy}{dx}\] by the chain rule.

OpenStudy (anonymous):

product rule??already,,,but dy/dx got -y/x,..then dont know..coz the answer 2y/x^2..

OpenStudy (anonymous):

You have to take the derivative again to find \[\frac{d^2y}{d^2x^2}\]

OpenStudy (anonymous):

yup,,but my answer didnt same like the ans given i got (-3x^2y^3) / (3x^3y^2)...can it be simplified???

OpenStudy (anonymous):

One sec. That doesn't seem right.

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

\[x^3\frac{d}{dx}[y^3] + y^3\frac{d}{dx}[x^3]\] \[=2x^3y^2\frac{dy}{dx} + 2x^2y^3 = 0\]\[\implies \frac{dy}{dx} = -\frac{y}{x}\]

OpenStudy (anonymous):

With me so far?

OpenStudy (anonymous):

yup...so far same..

OpenStudy (anonymous):

\[\implies \frac{d^2y}{d^2x^2} = -(\frac{1}{x}\frac{dy}{dx} + y\frac{d}{dx}[\frac{1}{x}])\] \[=-(\frac{1}{x}\cdot\frac{-y}{x} -\frac{y}{x^2})\] \[=\frac{2y}{x^2}\]

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

cant get it:((i use quotient rule..

OpenStudy (anonymous):

What? where?

OpenStudy (anonymous):

I used the product rule.

OpenStudy (anonymous):

\[\frac{dy}{dx} = -\frac{y}{x} = -yx^{-1}\] \[\]

OpenStudy (anonymous):

\[\implies \frac{d^2y}{d^2x^2} = \frac{d}{dx}[-yx^-1]\]\[=-y\frac{d}{dx}[x^{-1}] + x^{-1}\frac{d}{dx}[-y]\]\[= \frac{y}{x^2} + x^{-1}(-\frac{dy}{dx})\]\[ = \frac{y}{x^2} + x^{-1}(-(-yx^{-1}))\]\[= \frac{y}{x^2} + yx^{-2} \]\[=\frac{y}{x^2} + \frac{y}{x^2} \]\[=\frac{2y}{x^2}\]

OpenStudy (anonymous):

is another way if you prefer.

OpenStudy (anonymous):

ok2..i got it finally:) use quotient rule for -y/x then - ((x(dy/dx)-y) / x^2) then i subtitute dy/dx with -y/x...so got sme ans with u...

OpenStudy (anonymous):

yep. Should be the same either way. I just usually ignore quotient by using power rule with negative exponents and the product rule..

OpenStudy (anonymous):

then can u help me with this d/dx \[d/dx \left( 1/2 \tan^{-1} x + 1/4 \ln \left( x+1 \right)^{2}\div \left( x ^{2}+1 \right) \right)\]